OpenStudy (anonymous):

which of these is the equation of the line y-intercept 4 and x-intercept 4

thomaster (thomaster):

Where is the "these" part?

OpenStudy (anonymous):

A.) y=-x+4 B.)y=-4x C.) y=c-4 D.)4y=-4x

OpenStudy (anonymous):

A. You can say, for the y intercept we make x=0 and solve for y thus giving you 4 and for the x intercept, we say let y = 0 and thus find x = 4.

OpenStudy (anonymous):

remember the standard equation for a straight line is given by the expression: \[y=mx+c\]where c is the y intercept (i.e. letting x = 0 will give us y = c) and m is the gradient (steepness of the slope) given by the expression: \[m=\Delta y / \Delta x = (y _{2}-y _{1})/(x _{2}-x_{1})\]Hope this helps :)